OpenStudy (anonymous):

use the exponent rule to simplify the expression. assume all variables represent nonzero numbers. (m^4n)^-4/m^-8n^3 I got 1/m^8n is this correct?

OpenStudy (unklerhaukus):

\[\frac{(m^4n)^{-4}}{m^{-8}n^3}=m^{4\times-4+8}n^{-4-3} =m^{-8}n^{-7}=\frac 1{m^8n^7}\]